Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 17:02 Thursday 04 February 2021 (15928) Print this report
Oplev installation and alignment recovery.

With Hirata-san

As reported in a previous entry we installed the oplev. After the venting last year, the breadboards with the optics were removed, but their position on the breadboard and orientations did not change. Because of this I was expecting a quick recovery of the oplev but this did not happen. The highlights are:

  • I did not want to move the oplev optics omn the breadboards because they had been aligned in the past.
  • The collimator seemed to be too far off to the right of the board and the tip and tilt of the holder had to be adjusted by suspiciouly large amounts in order to get the beam onto the SR3 mirror. Initially it was going underneath the lower breadboard and onto the right of the central feature.
  • For some reason IM-Y was -2881 urad, I used the picomotor to bring it to zero.
  • I used the IP horizontal fishing rods to move the IP close to the setpoint.
  • We lost the original yaw of the chain when we reinstalled the F0 yaw motor with limit switches. We installed it in a way such that it should be close to the original orientation but for the purposes of the oplev it might be still far. We will need to align the mirror again using green light, for example.
  • The GAS filters were out of place and the vertical DoF couple with pitch and yaw likely through cables. Nevertheless, the oplev beam is way off to such a extent that this is likely not the problem.
  • Temperature: 23.1 degrees.

Given the above, the plan will be:

  • Set IM-P and IM-Y to the values we had in ALIGNED state.
  • Set the IP close to the setpoint.
  • Set the GAS filters close to the setpoint also. In the case of F0 adjust the ballast mass because too much actuation may be required.
  • Then move the collimator tombstone first and, second, adjust the tip and tilt of the first folding mirror.

The readout of the OSEMs during O3GK in ALIGNED state were:

  • IM-L: 204 um,
  • IM-T: 41 um,
  • IM-V: 174 um
  • IM-R: -205 urad
  • IM-P: -338 urad
  • IM-Y: 690 urad.
  • Guardian ALIGNED state: 500

The amount of actuation that would bring the GAS filters to their setpoints are

  • F1: -16,000 cnt,
  • BF: -4,500 cnt.
Images attached to this report
Comments to this report:
kouseki.miyo - 7:44 Friday 05 February 2021 (15935) Print this report

For people work in SR3, previous entry Fabian mentioned is klog#15921.

fabian.arellano - 16:21 Monday 08 February 2021 (15953) Print this report

With Washimi-san.

Pictures at SR3 remedying work after O3.

Alignment recovery

  • The IP and F0 are locked since last week. We don't need them unlocked for the adjustment of the ballast masses.
  • I moved IM-P from -640 urad to -342 urad (the aim was -339 urad per the commented klog entry)
  • I moved IM-Y to approximately zero. We will use the F0 yaw motor to adjust the mirror yaw.
  • We moved the position in which F0 is locked to -834 um, which is the current setpoint.
  • With certain amount of coil-magnet actuation we set BF and F1 close to their setpoints.
  • We moved the oplev collimator a few millimeters towards the -X direction.
  • Then, by adjusting the tip and tilt of the collimator the beam reached the tilt QPD but not to the center yet.
  • For the length sensing beam, we adjusted the lens holder sideways from the base to make the beam go through the centre of the lens; we used a ground glass alignment disk. See this picture.
  • Then we tilted the second folding mirror to make the beam reach the centre of the length sensing QPD.
  • Finally we adjusted the yaw of the suspension using the F0 yaw motor to bring the beam to the center of the tilt QPD.
  • The temperature was 22.8 degrees, which is a bit low. See entry 15949. Heaters were added. During our work we used the Koach filters and the temperature raised to 23 degrees. After our work it went down back to 22.8 degrees.

I measured transfer functions in IM-P, IM-Y and IM-R and the system looks healthy.

As preparation to the adjustment of the IM ballast mass I checked the BF fishing rod is not jammed, and I left it at the upper limit of its moving range.

Images attached to this comment
Search Help